# The Unbounded Denominators Conjecture

### v2

@@ 4.2. Modular Forms
for all $N$. The key leveraging step is contained in the following theorem:
Theorem 4.2.6.
Suppose that there exists an $N$ such that $R_{N}:M_{N}]>1$. Then, for every prime $p$ not dividing $N$, one has
$R_{Np}:M_{Np}]\geq 2[R_{N}:M_{N}].$
The proof of this theorem is in the next section 4.3 below. However, we now explain why Theorem 4.2.6 together with a sufficiently good bound on $R_{N}:M_{2}]$ implies the unbounded denominators conjecture.
@@
(4.2.8) $R_{N\prod_{p\in S}p}:M_{N\prod_{p\in S}p}]\geq[R_{N}:M_{N}]2^{\#S}>2^{(1-\varepsilon)X/\log X},$
by the prime number theorem. This quantity certainly increases faster than any power of $X$. On the other hand, from the assumed bound on $R_{N}:M_{2}]$ together with the bound (4.2.5), we obtain
@@
4.3. The proof of Theorem 4.2.6
We now put ourselves in the situation of Theorem 4.2.6. That is, we have $R_{N}$ is strictly larger than $M_{N}$. Recall that forms in $R_{N}$ are all invariant by a subgroup $G=G_{N}\subset\langle E,\Gamma(N)\rangle$ with $L(G_{N})=N$ by Lemma 4.2.3. Let $p$ be a prime with $(N,p)=1$.
The main idea of this section is to exploit the fact that if $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{N}$, then $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/Np}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since the form $f(\tau)$ is invariant under $G$, the form $f(\tau/p)$ is invariant under $AGA^{-1}$ and thus also the group $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $Np$. In particular $f(\tau/p)$ has cusp width dividing $Np$ at each cusp, and hence $f(\tau/p)\in R_{Np}$. Let $R_{N}M_{Np}$ denote the field generated by forms in $R_{N}$ and $M_{Np}$, which is contained in $R_{Np}$. We have
because the intersection of $M_{Np}$ and $R_{N}$ is $M_{N}$. Thus
$\frac{[R_{Np}:M_{Np}]}{[R_{N}:M_{N}]}=[R_{Np}:R_{N}M_{Np}]$
is an integer, and it is either $\geq 2$, in which case we have proven Theorem 4.2.6, or $R_{Np}=R_{N}M_{Np}$. In particular, in order to prove Theorem 4.2.6, we may assume that $R_{Np}$ is generated by $R_{N}$ and $M_{Np}$, and in particular that $f(\tau/p)$ is invariant under the group $\langle E,G\cap\Gamma(Np)\rangle$. This implies that $f(\tau/p)$ is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. That is, Theorem 4.2.6 is reduced to the following claim:
Proposition 4.3.1.
If $f(\tau)$ is invariant by a finite index sugroup $G$ containing $E$ with $L(G)=N$ and $f(\tau/p)$ it is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$, then $f(\tau)$ is invariant under a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$.
The remainder of this section is devoted to the proof of this proposition. We will exploit an idea due to Serre [Tho89] that shows $G$ and $AGA^{-1}$ are in some sense “far apart” when $G$ is not a congruence subgroup.
We first prove a preliminary lemma. Recall that $\Gamma_{0}(p)$ and $\Gamma^{0}(p)$ are the subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$ consisting of matrices congruent to $\left(\begin{matrix}*&*\\ 0&*\end{matrix}\right)\bmod p$ and $\left(\begin{matrix}*&0\\ *&*\end{matrix}\right)\bmod p$ respectively. There is an identification $A\Gamma_{0}(p)A^{-1}=\Gamma^{0}(p)$ inside $\mathrm{SL}_{2}(\mathbf{Z})$.
Lemma 4.3.2.
Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup with $L(G)=N$ in the sense of Definition 4.1.1. Let $H$ be the group generated by $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ together with $G\cap\Gamma(Np)$. Then $H$ contains $G\cap\Gamma(N)\cap\Gamma^{0}(p)$.
Proof.
Since $G$ and $H$ are both subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$, there are maps
(4.3.3) $\psi_{G},\psi_{H}:G,H\rightarrow\mathrm{SL}_{2}(\mathbf{Z})\rightarrow\mathrm{SL}_{2}(\mathbf{Z}/Np).$
The kernel of $\psi_{G}$ is equal to $G\cap\Gamma(Np)$, which is contained in $H$ by definition. Hence to show that $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ is contained in $H$, it suffices to show that the image of $\psi_{H}$ contains the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$. Since $G$ has finite index inside $\mathrm{SL}_{2}(\mathbf{Z})$ and Wohlfahrt level $N$, it contains the matrices
$\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right),\quad\left(\begin{matrix}1&0\\ N&1\end{matrix}\right).$
But it is easy to see that these two matrices generate the projection of $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ onto $\mathrm{SL}_{2}(\mathbf{F}_{p})$. Hence the image of $G\cap\Gamma(N)$ under $\psi_{G}$ is precisely the matrices $\gamma=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\mod Np$ congruent to $I\bmod N$, and the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ is precisely the matrices of the form
(4.3.4) $\left(\begin{matrix}a&bp\\ c&d\end{matrix}\right)\mod Np$
inside $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ which are congruent to $I\mod N$. Our task is to show that such matrices also lies the image of $H$. Since $(N,p)=1$, it follows that
(4.3.5) $b\equiv c\equiv 0\bmod N,\qquad a\equiv d\equiv 1\bmod N.$
We observe that the matrix
(4.3.6) $\left(\begin{matrix}a&b\\ cp&d\end{matrix}\right)\mod Np$
also lies in $\mathrm{SL}_{2}(\mathbf{Z}/Np)$; the determinant is unchanged, and the matrix is congruent to the identity modulo $N$ by equation (4.3.5). Thus it lies in the image of $G\cap\Gamma(N)$. Let $\gamma$ denote any lift to $G$; clearly the lower left entry is divisible by $p$. Hence
$A\gamma A^{-1}=\left(\begin{matrix}a&bp\\ c&d\end{matrix}\right)\in AGA^{-1}$
lies in $\mathrm{SL}_{2}(\mathbf{Z})$, and thus in $H$. This shows that the image of $H$ contains the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ and we are done. ∎
Since $G$ is normal and is contained in $\langle E,\Gamma(N)\rangle$, we may define the group $S$ by taking $S=\langle E,\Gamma(N)\rangle/G$. By construction, the group $S$ is finite. There is a natural projection:
@@
This is more or less precisely [Tho89, Theorem 3] with the addition of level structure as in [Ber94].
@@
Note that

### v3

@@ 4.3. A leveraging argument
for all $N$. We can now compare this bound against the one coming from Theorem 4.3.2
@@
(4.3.6) $R_{N\prod_{p\in S}p}:M_{N\prod_{p\in S}p}]\geq 2^{\#S}>2^{(1-\varepsilon)X/\log X},$
by the prime number theorem. The RHS certainly increases faster than any power of $X$. On the other hand, from the assumed bound on $R_{N}:M_{2}]$ together with the bound (4.3.4), we obtain
@@
4.4. Amalgams and a non-abelian version of Ihara’s Lemma
Since $G=G_{N}$ is normal and is contained in $\langle E,\Gamma(N)\rangle$, we may define a group $S$ by taking $S=\langle E,\Gamma(N)\rangle/G$. By construction, the group $S$ is finite. There is a natural projection:
@@ 4.4. Amalgams and a non-abelian version of Ihara’s Lemma
This is more or less precisely [Tho89, Theorem 3] with the addition of level structure as in [Ber94]. One may think of Lemma 4.4.1 as a non-abelian version of Ihara’s Lemma, because (as explained below in the proof of Lemma 4.6.2) the case when $S$ is a vector space over $\mathbf{F}_{q}$ reduces to the statement of Ihara’s Lemma as proved by Ribet [Rib84]. (The proofs of both claims are very similar.)
@@
Let $B\subset\mathrm{SL}_{2}(\mathbf{F}_{p})$ denote the Borel subgroup of upper triangular matrices. There is a natural surjection $\pi:\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow B$ whose kernel is $\Gamma(Np)$. We have the following extension of Lemma 4.4.1.
Lemma 4.4.4.
The map $(f_{1},f_{2},\pi):\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow S\times S\times B$ is surjective.
Proof.
Let $\gamma=\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right)$ and $\eta=\left(\begin{matrix}1&0\\ N&1\end{matrix}\right)$. The assumption that $L(G)=N$ and $G$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$ impies that $\gamma,\eta\in G$. Since $A^{-1}\gamma^{p}A^{-1}=\gamma\in A^{-1}GA\cap\Gamma_{0}(p)$, we see that $\gamma\in\ker(f_{1})$ and $\gamma\in\ker(f_{2})$, and yet
$\pi(\gamma)=\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right)\in B$
generates the normal unipotent subgroup $U\subset B$. By Goursat’s Lemma, we can detect the failure of surjectivity coming from a map of $S\times S$ and $B$ to some common quotient. Because the image contains $0\times 0\times\langle U\rangle$, this common quotient is a quotient of the abelian group $B/\langle U\rangle$. Thus, by Nakayama’s Lemma, the failure of surjectivity can be detected by maps to $\mathbf{F}_{q}$ for primes $q$. Maps to $\mathbf{F}_{q}$ are determined by cohomology classes with coefficients in $\mathbf{F}_{q}$. Let $\mathrm{S}G:=G\cap\Gamma(N)$. Since $E\in G$, we have
$S\simeq\langle E,\Gamma(N)\rangle/G\simeq\Gamma(N)/\mathrm{S}G,$
and so the map $(f_{1},f_{2})$ remains surjective after restriction to $\Gamma(N)\cap\Gamma_{0}(p)$. The surjectivity of $(f_{1},f_{2})$ implies the injectivity of the map
(4.4.5) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{\oplus 2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}),$
The assumption that $L(G)=N$ implies that
(4.4.6) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})\cap H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})=0\in H^{1}(\Gamma(N),\mathbf{F}_{q}),$
where $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})\subset H^{1}(\Gamma(N),\mathbf{F}_{q})$ denotes the classes which vanish after restriction to a congruence subgroup (Definition 4.5.1). This is because the kernel of any non-trivial map in $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})$ has level strictly divisible by $N$. The claim (4.4.5) follows from (4.4.6) as a consequence of Ihara’s Lemma, as proved by Ribet [Rib84] (see Lemma 4.6.2). The maps $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow B/\langle U\rangle\rightarrow\mathbf{F}_{q}$ on the other hand come from the classes in $H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q})$ which restricts to zero on $H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q})$, and thus what is required is to upgrade the injection of (4.4.5) to an injection
(4.4.7) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{\oplus 2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q}),$
which is dual to the desired claim that the map
$\Gamma(N)\cap\Gamma_{0}(p)\rightarrow S^{\mathrm{ab}}/qS^{\mathrm{ab}}\times S^{\mathrm{ab}}/qS^{\mathrm{ab}}\times B/U$
is surjective. But now we may invoke an enhanced version of Ihara’s Lemma (Lemma 4.6.3) which we prove in §4.6, and the injectivity of (4.4.7) follows directly from (4.4.6). ∎
Remark 4.4.8.
Because $\gamma$ and $E$ map to zero in $S\times S$ and $B/\langle E,U\rangle$ has order (p1)/2(p-1)/2, The proof of Lemma 4.4.4 is almost immediate if one imposes the additional hypothesis that $\left(\frac{p-1}{2},|S|\right)=1$. In particular, one would not have to appeal to the results in §4.5 and §4.6 (which are not used elsewhere in this paper). It turns out that proving Lemma 4.4.4 under this weaker hypothesis would suffice for the proof of the unbounded denominators conjecture. The key point is that if $\langle E,\Gamma(N)\rangle/G_{N}\simeq S$ and $G_{Np}$ is the group given by the intersection of the three groups $\langle E,\Gamma(Np)\rangle$, $G$, and $AGA^{-1}$, then $\langle E,\Gamma(Np)\rangle/G_{Np}\simeq S\times S$. In particular, one can control the primes dividing $S$ as one varies $N$. Then, in the argument of Proposition 4.3.5, instead of adding all primes <X<X prime to $N$, one only includes primes in some arithmetic progression satisfying the congruence $\left(\frac{p-1}{2},|S|\right)=1$ for some fixed $S$. However, it seems more natural to prove Lemma 4.4.4 without such an ugly hypothesis. Additionally, sections §4.5 and §4.6 may be of independent interest.
Returning to the assumptions of Lemma 4.4.4, let $K=\langle\ker((f_{1},\pi),\ker(f_{2})\rangle$ be the group generated by $\ker((f_{1},\pi))$ and $\ker(f_{2})$. We deduce from Lemma 4.4.4 that the image of $(f_{1},f_{2},\pi)$ contains the elements $(x,0,z)$ and $(0,y,0)$ for any triple $(x,y,z)\in S\times S\times B$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker((f_{1},\pi))$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that $K=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, or equivalently that
(4.4.9) $\langle E,G\cap\Gamma(Np),A^{-1}GA\cap\Gamma_{0}(p)\rangle=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p).$
We now complete the proof of Theorem 4.3.1 and hence the proof of Theorem 4.3.2 (as explained at the beginning of § 4.3).
Proof of Theorem 4.3.1 .
Consider the function $f(p\tau)$. Assume that this lies in the algebra generated by $f(\tau)$ and $M_{Np}$. Then $f(p\tau)$ is invariant under both $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. But from (4.4.9) we see that these groups together generate a congruence subgroup, and thus $f(p\tau)$ and $f(\tau)$ are congruence, a contradiction. ∎
4.5. Invariant vectors
We recall the following definition (cf. [CV19, §3.7]).
Definition 4.5.1.
Let $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a congruence subgroup. A congruence class $\eta\in H^{1}(\Gamma,\mathbf{F}_{\ell})$ is a class that restricts to zero on some congruence subgroup $\Gamma^{\prime}\subset\Gamma$. Denote the subgroup of congruence classes by
$H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{\ell})\subset H^{1}(\Gamma,\mathbf{F}_{\ell}).$
If $\widehat{\Gamma}$ denotes the congruence completion of the group $\Gamma$, then $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{\ell})\simeq H^{1}(\widehat{\Gamma},\mathbf{F}_{\ell})$. For a prime $\ell$, one may define ([CE12]) the groups
$\widetilde{H}^{1}(\mathbf{F}_{\ell}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{F}_{\ell}),\quad\widetilde{H}^{1}(\mathbf{Q}/\mathbf{Z}):=\lim_{N}H^{1}(\Gamma(N),\mathbf{Q}/\mathbf{Z}).$
over all levels $N$. The limit has an action of the group $\mathrm{SL}_{2}(\widehat{\mathbf{Z}})=\prod_{p}\mathrm{SL}_{2}(\mathbf{Z}_{p})$. The goal of this section is to prove:
Theorem 4.5.2.
[…]
Corollary 4.5.3.
Let $N$ be an integer,and $\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$. If, for all $g\in\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$, the class $g\eta-\eta\in H^{1}(\Gamma(N),\mathbf{F}_{\ell})$ is a congruence, then $\eta$ is congruence.
Proof.
[…]
Our first goal is to control the group $H^{2}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})$ for various $N$, in particular for $N=1$, which we do in a sequence of steps.
Lemma 4.5.4.
We have $H_{2}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{Z})=0$ for all primes $p$.
Proof.
It suffices to prove the vanishing of $H_{2}(\Delta,\mathbf{Z})$ for any Sylow subgroup of $\mathrm{SL}_{2}(\mathbf{F}_{p})$. For odd primes, the Sylow subgroup is cyclic and the cohomology of a cyclic group is only non-zero in even degree. The $2$-Sylow subgroup is a generalized quaternion group, whose cohomology also vanishes in odd degree (as follows from [Hup67, Satz 25.3(a), p.643] and [Swa60, Theorem 2]). ∎
Lemma 4.5.5.
For $n=1$ and $n=2$, we have:
$H^{n}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{\ell})^{\vee}\simeq H_{n}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{\ell})=\begin{cases}\mathbf{F}_{\ell},&p=\ell\in\{2,3\}\\ 0,&\text{otherwise}.\end{cases}$
Proof.
There is a short exact sequence:
$0\rightarrow H_{2}(G,\mathbf{Z})/\ell\rightarrow H_{2}(G,\mathbf{F}_{\ell})\rightarrow H_{1}(G,\mathbf{Z})[\ell]\rightarrow 0,$
and $H_{1}(G,\mathbf{F}_{\ell})\simeq H_{1}(G,\mathbf{Z})/\ell$. Hence the result follows from combining Lemma 4.5.4 with the fact that $\mathrm{SL}_{2}(\mathbf{F}_{p})^{\mathrm{ab}}$ is trivial for $p\geq 5$ and $\mathbf{Z}/p\mathbf{Z}$ for $p=2$ and $p=3$. ∎
Lemma 4.5.6.
For $n=1$, we have:
$H^{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})=\begin{cases}\mathbf{F}_{\ell},&p=\ell\in\{2,3\}\\ 0,&\text{otherwise}.\end{cases}$
For $n=2$, we have $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})=0$ unless $\ell=p$ and $p\leq 5$.
Remark 4.5.7.
We shall compute the exceptional cases when $\ell=p\leq 5$ in Lemma 4.5.11 below as a consequence of Theorem 4.5.2.
Proof.
Assume that $\ell\neq p$. By Hochschild–Serre, we have an isomorphism
$H^{*}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})\simeq H^{*}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{\ell})$
and thus the result follows from Lemma 4.5.5. Thus we may assume that $\ell=p$. Assume that $p>2$. Let $G(p)$ be the $p$-congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z}_{p})$. We may assume that $p>2$. Then $G(p)$ is $p$-torsion free and $p$-powerful, so, with $M=M_{0}(\mathbf{F}_{p})^{\vee}$ where $G(p)/G(p^{2})\simeq M_{0}(\mathbf{F}_{p})$, we deduce by Lazard’s Theorem ([Laz65, Chapter V, 2.2.6.3 and 2.2.7.2, page 167]) that there are isomorphisms
$H^{1}(G(p),\mathbf{F}_{p})\simeq M,\quad H^{2}(G(p),\mathbf{F}_{p})\simeq\wedge^{2}M,$
where the cup product map $\wedge^{2}M:H^{1}\wedge H^{1}\rightarrow H^{2}$ is an isomorphism. Assuming $p\geq 3$, we find that $M\simeq M^{\vee}$ is self-dual as a $\mathrm{SL}_{2}(\mathbf{F}_{p})$-module and so $\wedge^{2}M\simeq M$. Moreover, we have  $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$. Consider the Hochschild–Serre spectral sequence:
$H^{i}(\mathrm{SL}_{2}(\mathbf{F}_{p}),H^{j}(G(p),\mathbf{F}_{p}))\Rightarrow H^{i+j}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p}).$
Since $H^{1}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{p})=H^{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})$ and $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$, there is an exact sequence:
(4.5.8) $0\rightarrow H^{2}(\mathrm{SL}_{2}(\mathbf{F}_{p}),\mathbf{F}_{p})\rightarrow H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})\rightarrow H^{1}(\mathrm{SL}_{2}(\mathbf{F}_{p}),M).$
If $p\neq 5$, then $H^{1}(\mathrm{SL}_{2}(\mathbf{F}_{p}),M)=0$ (see [DDT97, Lemma 2.48]) and the result follows from Lemma 4.5.5. ∎
We deduce:
Lemma 4.5.9.
For every prime $\ell$, there is an isomorphism $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\simeq H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{\ell}),\mathbf{F}_{\ell})$. If $N$ is a power of $\ell$ and $G(N)\subset\mathrm{SL}_{2}(\mathbf{Z}_{\ell})$ the corresponding principal congruence subgroup, then
$H^{2}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})\simeq H^{2}(G(N),\mathbf{F}_{\ell}).$
If $\ell$ is odd and $N$ is a non-trivial power of $\ell$ or $N\geq 8$ is a power of $\ell=2$, then the map
$H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\rightarrow H^{2}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})$
is trivial.
Proof.
Since $H^{n}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})=0$ for $n=1$ and $n=2$ unless $\ell=p$, the first two claims follow from the Künneth formula and Lemma 4.5.6. It remains to show that the map
$H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})\rightarrow H^{2}(G(N),\mathbf{F}_{p})$
is the zero map for $N=p$ if $p$ is odd and $N=8$ if $p=2$. For $p>2$, we have $H^{2}(G(N),\mathbf{F}_{p})\simeq M$ and $M^{\mathrm{SL}_{2}(\mathbf{F}_{p})}=0$. Since the source is $\mathrm{SL}_{2}(\mathbf{F}_{p})$-invariant, the image must be trivial. For $p=2$, we have $H^{2}(G(N),\mathbf{F}_{2})\simeq\wedge^{2}(M)\simeq M$ whenever $N\geq 4$ (to ensure that $G(N)$ is $2$-powerful). Unlike what happens for $p$ odd, we have $M^{\mathrm{SL}_{2}(\mathbf{F}_{2})}=\mathbf{F}_{2}$. However, the map
$M=H^{1}(G(4),\mathbf{F}_{2})\rightarrow H^{1}(G(8),\mathbf{F}_{2})=M$
is zero, and thus the induced map
$\wedge^{2}M=H^{2}(G(4),\mathbf{F}_{2})\rightarrow H^{2}(G(8),\mathbf{F}_{2})=\wedge^{2}M$
is also zero. ∎
Now let us consider the following commutative diagram for $N\in\{3,4,5,8,16\}$ and $\ell$ dividing $N$ coming from compatible Hochschild–Serre spectral sequences:
[…]
Here $H^{2}(\Gamma(N),\mathbf{F}_{\ell})=0$ because $\Gamma(N)$ is a free group. The last vertical map is zero by the previous lemma if $N\in\{3,5,8,16\}$, and thus the image of $(\widetilde{H}^{1})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$ in $(\widetilde{H}^{1})^{\widehat{\Gamma}(N)}$ lands in the image of $H^{1}(\widehat{\Gamma}(N),\mathbf{F}_{\ell})$ in these cases. But these are finite groups we can compute explicitly.
Lemma 4.5.10.
We have
$\dim H^{1}(\Gamma(3),\mathbf{F}_{3})^{\mathrm{SL}_{2}(\mathbf{F}_{3})}=0,$
$\dim H^{1}(\Gamma(5),\mathbf{F}_{3})^{\mathrm{SL}_{2}(\mathbf{F}_{5})}=0,$
$\dim H^{1}(\Gamma(4),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/4\mathbf{Z})}=\dim H^{1}(\Gamma(8),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})}=\dim H^{1}(\Gamma(16),\mathbf{F}_{2})^{\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})}=1.$
For $N=3,4,5$, the same result holds even after considering the semi-simplifications of these modules.
Proof.
Recall that for $N=3$, 44, and $5$ that $X(N)$ has genus zero. Hence the cohomology $V=H^{1}(\Gamma(N),\mathbf{Z})$ is coming entirely from the the cusps, which correspond to the cosets of $\langle\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\rangle$ in $\mathrm{PSL}_{2}(\mathbf{Z}/N\mathbf{Z})$. In particular, in the Grothendieck group,
$V_{\mathbf{Q}}:=\left[H^{1}(\Gamma(N),\mathbf{Q})\right]\simeq\mathbf{Q}\left[\mathrm{PSL}_{2}(\mathbf{Z}/N\mathbf{Z})/\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\right]-[\mathbf{Q}].$
Since $\Gamma(N)$ is free, this is enough to determine the semi-simplification of $V_{\mathbf{F}_{\ell}}:=H^{1}(\Gamma(N),\mathbf{F}_{\ell})$. We consider each case in turn.